A women has $56,000. Part is invested for 1year in a CD paying 7% interest, and the remaining amount in a cash managements fund paying 10%. The total interest on the two investments is $4,220. How much does she invest at each rate?
Answer by mananth(16946)

A women has $56,000. Part is invested for 1year in a CD paying 7% interest, and the remaining amount in a cash managements fund paying 10%. The total interest on the two investments is $4,220. How much does she invest at each rate?
...
let x be invested in 7%
56000-x = invested in 10%
...
interest = 4220
..
0.07x+(56000-x)*0.1= 4220
0.07x +5600-0.1x=4220
-0.03x=4220-5600
-0.03x=-1380
x= -1380/0.03
x= $46,000 @ 7%
Balance at 10%
m.ananth@hotmail.ca